Operations Report – January 4th

Howdy Mission Support!

I am about to run out to the ScienceDome to check the SOC

Crew 202 Operations Report 04-JAN-2019

SOL: 5

Name of person filing report: Kasey Hilton

Non-nominal systems: None

Notes on non-nominal systems: Nothing to report

Generator (hours run): 16hr 29min; Turned on last night (03Jan2019) at 16:36; Turned off this morning (04Jan2019) at 9:05; Turned on tonight (04Jan2019) at 15:02

Solar SOC – Turned on (03Jan2019) 83%; Turned off (04Jan2019) 100%; Turned on (04Jan2019) 81%

Diesel Reading – 73%

Propane Reading – 48%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – About 75%; 400 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 75%; 230 gallons (assuming that the tank is 300 gallons)

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 18:06 to refill tank

Water Meter: 01398024 units

Toilet tank emptied: Yes, still very stinky ://

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Yes

Hours: 64.7

Beginning charge: 100%

Ending charge: 44%

Currently charging: Yes

Opportunity rover used: No; Still out of brake fluid

Hours: 45.4

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 62.9

Beginning charge: 100%

Ending charge: 61%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: Not applicable

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Not used

CrewCar used and why, where? Off site

General notes and comments: Today was the first day that nothing had broken!

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Best,

Kasey